HOW TO PROOF THIS QUESTION
 Q. COS C = -COS(A+B).
PLEASE TELL ME
 

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
As we know
 
in a triangle
 
A+ B + C = 180
 
Now
 
A+B = 180 – C
 
take cos both sides
 
cos (A+ B) = cos (180 – C)
 
cos A+B = – cos C
 
cos C = – cos (A+B)